Estimate the number of electrons in 100g of water. How much is the total negative charge on these electrons ?

Grade:

3 Answers

Arun
5 years ago
N=100/18=> no. Of molecule`s=[(100/18*6.023*10²³)]One molecule of water has (2+8)electrons Ne-=[100/18*6.023*10²³]Total charge=1.6*10-¹9 *Ne-v
Gitanjali Rout
5 years ago
=100/18=> no. Of molecule`s=[(100/18*6.023*10²³)]One molecule of water has (2+8)electrons Ne-=[100/18*6.023*10²³]Total charge=1.6*10-¹9 *Ne-vmb
Gitanjali Rout
5 years ago
hjnb h  ghvj kj.........................................................................................................................................................................................................